OpenStudy (anonymous):

How do you multiply and simplify 1/35*400d

OpenStudy (anonymous):

make 400d/1 and multiply across \[\left(\begin{matrix}1 \\ 35\end{matrix}\right) * \left(\begin{matrix}400d \\ 1\end{matrix}\right)\]\[\rightarrow \left(\begin{matrix}400d \\ 35\end{matrix}\right)\rightarrow \left(\begin{matrix}80d \\ 7\end{matrix}\right)\]